# Redistricting Reforms Reduce Gerrymandering by Constraining Partisan Actors

### v2

@@ 3.3 Equilibrium solution as a treatment variable
As might be expected, states with independent redistricting commissions, such as the running example of Michigan, have a realized and maximum leeway of zero. In 2010, Michigan redistricting was entirely controlled by the legislature and governor, who were both Republican. By 2020, the process was reformed to have a strong independent commission. In the new process, the commission is bound by strict criteria, the process allows for court review, and even the commissioners themselves are selected in part by a lottery system. With the reform, Michigan’s maximum leeway changed from the highest observed value to zero, meaning that partisan actors have no control over redistricting outcomes, on average.
Other states’ scores are similarly interpretable, where introducing a commission (e.g., New York, Virginia, or Colorado) drastically reduces its leeway to match states with similar sets of rules in 2020. States that had intervening litigation to clarify the interpretation of state redistricting rules (e.g., New Mexico, North Carolina, or Pennsylvania) see appropriate changes in leeway, even absent a commission. Further, in our party-signed treatment, states that had a flip in party control of the whole system see large changes in leeway in the correct direction (e.g. West Virginia). Finally, states with minor changes to the total control of state institutions see minor changes in leeway (e.g., North Carolina).
3.4 Validation
@@
The model was designed to be flexible enough to capture the actual redistricting processes in each state. Any misspecification is therefore due to how the moves by nature are specified, or in the underlying setup of two opposing partisan actors competing on a single zero-sum dimension. First, we find that the treatment values are not sensitive to specific parameter values, up to monotonic transformations. Across the 100 random draws from the prior, the average pairwise Spearman correlation between the Nash equilibria for each state is above 0.99.888i.e., for each pair of random draws, we calculate the Spearman correlation between the two vectors of calculated state equilibria, then average these correlations across all pairings. This also gives us confidence that the specific choice of the prior is not influencing the results.
@@
Legislature 31.9 0.0 19.1 51
Commission 0.0 18.9 3.0 22
Court 1.7 0.0 12.3 14
Total 33.6 18.9 34.4 87
We also compare the equilibrium path in the model with the actual institution that drew the final map in each state. Because the moves by nature in the game are random, the actual equilibrium path may be a probabilistic mixture over multiple possible paths. We average these probabilities over the 100 draws of the game to arrive at an overall probability that a legislature, commission, or court draws the final map in each state, given its procedure and party control.
@@ 4.1 Difference-in-differences design
This assumption is analogous to the parallel trends assumption in the traditional binary difference-in-differences design, but differs in that it refers to changes in dosage across a continuous measure, rather than a single level change from zero. This conditional parallel trends assumption allows us to identify our estimand as
@@ 4.2 Adjusting for political geography
To make the adjustment, we use representative sets of nonpartisan redistricting plans within each state that respect each state’s specific redistricting rules. These simulation samples were separately generated for 2020 (McCartan et al.,, 2022) and 2010 (Kenny et al.,, 2024) using the algorithm of McCartan and Imai, (2023). We subtract the mean outcome in each state’s simulated sample, denoted by $\widetilde{Y}_{it}$, from the observed outcome $Y_{it}$ before estimating the causal effects under the DiD design. We argue that this subtraction of the outcomes based on simulated baseline plans from the observed outcomes accounts for the change in political geography in each state.
@@ 4.3 Estimation
In terms of the prior specification, the intercept receives a weakly informative prior of $\mathcal{N}\quantity(0,(2.5\cdot\sigma_{y})^{2})$, where $\sigma_{y}^{2}$ is the variance of $\Delta Y_{t}$. Each main effect coefficient receives a weakly informative prior of $\mathcal{N}\quantity(0,(0.75\cdot\sigma_{y}/\sigma_{x})^{2})$, where $\sigma_{x}^{2}$ is the variance of each covariate $x$. The prior standard deviation is further scaled by 0.25 for interaction terms, following the discussion in Gelman et al., (2008). Finally, the residual standard deviation receives an $\operatorname{Expo}(1/\sigma_{y})$ prior. Computation is carried out via NUTS-HMC as implemented in the Stan probabilistic programming language (Stan Development Team,, 2024).
@@ 5 Estimated Causal Effects of Redistricting Reforms
The outcome of the redistricting process is a complete congressional districting plan. A plan can be evaluated in a number of ways—how many seats each party is expected to win, how many of the seats are competitive, how well the partisan composition of the delegation matches the voters’ preferences, and so on. We begin this section by introducing quantitative measures of several partisan and nonpartisan aspects of districting plans. These constitute the outcome variables for the causal estimates we present in the second half of the section.
@@ 5.1 Outcome measures
Our primary partisan measure of districting plans is the expected number of seats won by the Republican party. As discussed in Section 4, we subtract the average outcomes in the simulated baseline redistricting plans (McCartan et al.,, 2022; Kenny et al.,, 2024) from the observed outcome in the enacted plan to adjust for the state-specific change in political geography, which is a potential confounding factor (Cottrell,, 2019). After this adjustment, the seat outcome ranges from $-1.57$ in Illinois in 2020 (favoring Democrats) to $1.78$ in Texas in 2020 (favoring Republicans). These simulation-differenced seats can be directly interpreted as a measurement of bias due to partisan gerrymandering, with positive values indicating a Republican bias beyond what would be expected based on the state’s political geography alone.
@@
We also study two more outcome measures based on the efficiency gap, which is a popular measure of partisan bias calculated as the difference between the wasted votes of the two parties (Stephanopoulos and McGhee,, 2015). Wasted votes are votes cast for losing candidates and votes cast for winning candidates beyond the threshold needed to win. The idea behind the efficiency gap is that a fair plan should have similar numbers of wasted votes for both parties. The partisan-signed version of the efficiency gap indicates which party benefits from the plan in terms of vote efficiency; we use positive values to indicate pro-Republican bias. Unlike with Republican seat share, we do not subtract the average simulated value from the partisan-signed efficiency gap measure. This is because the efficiency gap is already designed to have a universal scale, where 0 indicates partisan fairness; subtracting the simulated average would change the definition of a fair baseline for this measure. To create a nonpartisan measure, as with Republican seats, we take the absolute value of the difference between the enacted and average simulated plan’s efficiency gap. Smaller values indicate a fairer plan by this measure (closer to the normatively preferred value of 0).
@@ 5.2 Findings
Given the fitted outcome model, we can estimate the conditional average treatment effect (CATE) for any given starting and final value of maximum leeway and any combination of covariates. The right panel of Figure 5 plots these CATEs for each state that experienced a change in maximum leeway between 2010 and 2020, using the specific covariates of that state. There is a clear dose-response relationship between the decrease in maximum leeway and the decrease in the expected number of Republican seats. [^14] This pattern was also observed in robustness checks that used a flexible Bayesian Additive Regression Trees (BART) model for estimation. A state with Michigan’s dose (nearly 4) has an estimated CATE of around $-1.25$, with 95% credible interval (–2.0, –0.4), and so would be expected to gain just over 1 Democratic seat, on average. Similarly, a state with Ohio’s dose (just over 1) and covariates would be expected to gain about half of a Democratic seat (–0.6, –0.1). Motivated by this pattern, we also include in the right figure a dose-response curve calculated by calculating a CATE for dose level and averaging over the observed covariate distribution. [^15] Because of the interaction between the dose variable and the covariates (which do not have mean zero), the overall slope estimate and its variance differ from the estimate and variance of the coefficient on Dose in the linear model.
@@
In terms of partisan outcomes, the estimated ACRs on Republican seats are all positive, for both realized and maximum leeway. This is in line with an interpretation that increasing leeway benefits the party in control, as one might expect. Recall that the doses from Figure 3, for the realized leeway estimates, a dose of 4 represents approximately going from a Democratic legislature to an independent commission or an independent commission to a Republican legislature. For such a dose of 4, we would expect Republicans to gain around 0.7 seats (0.21, 1.18), or 7.2 simulation standard deviations (2.4, 12.2)—again, a substantial effect. We do not find a consistent effect on the signed efficiency gap outcome, which was not adjusted using the simulations.
@@ 6 Redistricting Commissions: Counterfactual Policy Analysis
What if redistricting reforms were adopted nationwide to constrain partisan actors? Redistricting commissions are the most commonly adopted map-drawing reform. A total of 15 states, including Michigan, New Jersey, and Arizona, currently use some form of commission. Advocates often argue that commissions can lead to fairer redistricting plans by further removing legislators from the process (e.g., American Academy of Arts & Sciences Commission on the Practice of Democratic Citizenship,, 2020). [^17] See also People Not Politicians Oregon or Ohio’s Citizens Not Politicians
In this section, we conduct a counterfactual policy analysis of the nationwide adoption of redistricting commissions. Though hypothetical, counterfactual policy analyses are useful tools for evaluating institutional reform proposals (Cervas and Grofman,, 2019). Our approach is well suited for this analysis, in part because commissions can vary drastically in the extent to which they limit the influence of partisan actors. We use our causal models to predict electoral outcomes under a series of hypothetical scenarios, in which every state adopts a redistricting commission with particular institutional powers. Although each of the 15 states uses a redistricting commission with different structures and rules, our model is able to characterize these different commissions in terms of their partisan leeway and use these treatment values to predict electoral outcomes under a given commission structure.
@@
We visualize the impact of each reform structure using a seats-votes curve. Seats-votes curves are often used to study a measure of partisan bias, operationalized as the difference in the share of seats and votes at a given point on the curve (Tufte,, 1973; King and Browning,, 1987; Katz et al.,, 2020). A seats-votes curve for which a 50% vote share translates exactly to a 50% seat share represents the baseline for partisan bias. We approximate the seats-votes curve around 50% vote share as a line, and estimate the effect of reforms on its slope and intercept separately. Up to a linear transformation, the slope corresponds to responsiveness and the intercept to seat share. These two outcome variables are differenced using the simulation sets, as discussed in Section 4.
We find that nationwide adoption of commissions would generally reduce partisan bias, and that reforms that place additional restrictions on partisan actors (as in Michigan) are generally more effective. Figure 7 shows a seats-votes curve for each counterfactual commission scenario (the orange lines), which display the predicted number of total Democratic seats in the US House (y-axis) for a given national Democratic vote share (x-axis). The model predicts that all commission structures would reduce the existing Republican advantage in the House driven by both the inefficient geographic distribution of Democratic voters (Chen and Cottrell,, 2016; Chen and Rodden,, 2015) and net partisan manipulation favoring Republicans (Kenny et al.,, 2023). We also compare these estimated relationships to the actual seats-votes curves for 2010 and 2020 (the black lines). We note that the lines for 2010 and 2020 intersect the 50% vote line below the 50% seat line, since a combination of the underlying geographic distribution of Democratic votes and gerrymandering efforts nationwide disadvantages the Democratic party.
@@
We find that New York- and Ohio-style reforms produce relatively moderate improvements in partisan bias and responsiveness. The point estimates for the number of Democratic seats gained under these reforms is 5.9 for New York-style reforms, with 95% credible interval (–0.39, 12.02), and 8.2 for Ohio-style reforms (1.1, 15.2). Responsiveness for New York- and Ohio-style reforms is also similar as for every 1pp increase in vote share, Democrats gain an estimated 8 seats with New York-style reforms, and 8.1 seats with Ohio-style reforms. These responsiveness values are only slightly higher than the actual responsiveness following 2020 redistricting. Ohio-style reforms appear more effective overall, though the substantive differences are small; we are more than 99% confident that Ohio-style reforms produce a larger effect on partisan bias than New York-style reforms and 86% confident that Ohio-style reforms increase responsiveness more than New York-style reforms do.
@@
Further, we find that Michigan-style reforms have the greatest effects in both responsiveness and Democratic seats, since partisan actors are the most constrained by the presence of a nonpartisan commission, no partisan veto points, and the potential for court review. We estimate that the number of Democratic seats gained under Michigan-style reforms is 22 with 95% credible interval (7.4, 38.3), and that for every 1pp increase in vote share, Democrats gain an estimated 9 seats (6.8, 11.2), an increase of 1.2 over the baseline. We are more than 99% confident that Michigan-style reforms have a larger effect than Ohio reforms, and more than 99% confident in a larger effect than New York-style reforms. We are also 87% confident that Michigan-style reforms increase responsiveness.
@@
Though all three proposed reforms reduce the deviation from partisan symmetry by increasing net Democratic seats, Michigan-style reforms suggest that much greater effects are produced when multiple nodes of the redistricting game constrain leeway by ensuring partisan nodes are nonpartisan, and do not precede partisan vetoes. Meanwhile, examples of New York-style and Ohio-style reforms demonstrate how constraining partisan actors at different nodes, through different reforms, may produce substantively similar effects.

### v3

@@ 3.4 Realized and maximum leeway scores
Intuitively, states introducing a commission (e.g., New York, Virginia, or Colorado) drastically reduced their leeway to match states with similar sets of rules in 2020. States that had intervening litigation to clarify the interpretation of state redistricting rules (e.g., New Mexico, North Carolina, or Pennsylvania) see appropriate changes in leeway, even absent a commission. Further, in our party-signed treatment, states that had a flip in party control of the whole system see large changes in leeway in the correct direction (e.g. West Virginia). Finally, states with minor changes to the total control of state institutions see minor changes in leeway (e.g., North Carolina).
3.5 Full process example: Michigan
@@
[…]
Our overall equilibrium solution for Michigan’s 2020 structure is a bias of 0.03-0.03, suggesting very little bias in favor of either party. This near-zero bias reflects the lack of partisan control in this particular structure. Because our solution approach reflects institutional designs at each time period, our equilibria estimates also capture intuitive changes in leeway over time. For example, in 2010, Michigan redistricting was entirely controlled by the legislature and governor, who were both Republican. This game structure results in the highest possible value of maximum leeway (party-blind) near 4, as a single party controls all important game steps. However, our near-zero 2020 equilibria reflects the adoption of a strong independent commission. In the new process, the commission is bound by strict criteria, the process allows for court review, and even the commissioners themselves are selected in part by a lottery system. With the reform, Michigan’s maximum leeway changed from the highest observed value to zero, meaning that partisan actors have no control over redistricting outcomes, on average.
3.6 Empirical validation of the proposed treatment vaiables
[…]
The model was designed to be flexible enough to capture the actual redistricting processes in each state. Any misspecification is therefore due to how the moves by nature are specified, or in the underlying setup of two opposing partisan actors competing on a single zero-sum dimension. First, we find that the treatment values are not sensitive to specific parameter values, up to monotonic transformations. Across the 200 random draws from the prior, the average pairwise Spearman correlation between the Nash equilibria for each state is above 0.99.888i.e., for each pair of random draws, we calculate the Spearman correlation between the two vectors of calculated state equilibria, then average these correlations across all pairings. This also gives us confidence that the specific choice of the prior is not influencing the results.
@@ 3.6 Empirical validation of the proposed treatment vaiables
Legislature 26.9 0.0 20.1 47
Commission 1.9 18.1 3.0 23
Court 5.8 0.9 10.3 17
Total 34.6 19.0 33.4 87
We also compare the equilibrium path in the model with the actual institution that drew the final map in each state. Because the moves by nature in the game are random, the actual equilibrium path may be a probabilistic mixture over multiple possible paths. We average these probabilities over the 200 draws of the game to arrive at an overall probability that a legislature, commission, or court draws the final map in each state, given its procedure and party control.
@@ 4 Estimating the Causal Effects of Institutional Leeway
We now discuss how our approach addresses the remaining challenges regarding outcome complexity and confounding bias in causal effect estimation (see the “estimation” and “outcome” boxes of Figure 1). We use a difference-in-differences design with the continuous treatment variable of institutional leeway to adjust for both observed and unmeasured state-specific confounding factors. We also use a simulation approach to adjust for changes in political geography that have happened during this time period. Specifically, we generate a representative sample of non-partisan redistricting plans for both 2010 and 2020 redistricting cycles to quantify state-specific changes in political geography. By accounting for such changes, we are able to better isolate the impact of redistricting reforms from that of political geography.
@@ 4.1 Difference-in-differences design
This assumption is analogous to the parallel trends assumption in the traditional binary difference-in-differences design, but differs in that it refers to changes in dosage across a continuous measure, rather than a single level change from zero. Compared to a traditional selection-on-observables assumption, the conditional parallel trends does not rule out confounding factors that are constant over time. With this assumption, we can identify our estimand as
@@
Evaluating the plausibility of the parallel trends assumption here is complicated by a lack of public data on precinct-level election returns for the 2000 redistricting cycle or earlier cycles. Without this data, we cannot generate simulated districts for decades before 2010, which prevents us from conducting placebo checks using the pre-trends, a common practice for differences-in-differences studies. We are able, however, to conduct a placebo check using an outcome that should be unaffected by redistricting reform, as we discuss in Section 5.3 below.
Thus, at least partially, we are left to justify the assumption based on substantive grounds. To violate the conditional parallel trends assumption, the change in outcomes (such as Democratic seats) that would have been observed in Michigan, had it not adopted reforms, would have to systematically differ from the corresponding change in other non-reform states like Wisconsin, after controlling for covariates. The covariates we include are some of the strongest predictors of treatment adoption and redistricting outcomes, which substantially increases the plausibility of this assumption.
Second, aspects of the treatment and outcome themselves weigh against the possibility of a violation. Changes in redistricting procedure are usually advanced outside of legislatures, often via citizen referenda funded by nonpartisan good-governance groups who are not responsible for map-drawing. The political parties also coordinate redistricting strategy at a national level. Further, state legislatures and other map-drawing bodies experience turnover between decades.
Together, these factors suggest that the change in Michigan’s outcomes would have looked much like that of a non-reform state’s, had it not adopted reforms. In the specific example of Wisconsin, political developments since 2023 further justify this conclusion. A narrow state supreme court victory by a justice who campaigned, in part, in opposition to gerrymandering, led to a successful court challenge to state legislative maps that resulted in a significant decrease in the partisan bias of those maps (on common bias metrics). Challenges to the congressional maps are currently pending.
@@ 4.2 Adjusting for political geography
[…]
@@ 4.3 Estimation of causal effects
[…]
@@ 5 Estimated Causal Effects of Redistricting Reforms
[…]
@@ 5.1 Outcome measures
Our primary partisan measure of districting plans is the expected number of seats won by the Republican party. As discussed in Section 4, we subtract the average outcomes in the simulated baseline redistricting plans (McCartan et al.,, 2022; Kenny et al.,, 2024) from the observed outcome in the enacted plan to adjust for the state-specific change in political geography, which is a potential confounding factor (Cottrell,, 2019). After this adjustment, the seat outcome ranges from 1.57-1.57 in Illinois in 2020 (favoring Democrats) to $1.78$ in Texas in 2020 (favoring Republicans). These simulation-differenced seats can be directly interpreted as a measurement of bias due to partisan gerrymandering, with positive values indicating a Republican bias beyond what would be expected based on the state’s political geography alone.
@@ 5.2 Empirical findings
Given the fitted outcome model, we can estimate the conditional average treatment effect (CATE) for any given starting and final value of maximum leeway and any combination of covariates. The right panel of Figure 6 plots these CATEs for each state that experienced a change in maximum leeway between 2010 and 2020, using the specific covariates of that state. There is a clear dose-response relationship between the decrease in maximum leeway and the decrease in the expected number of Republican seats. [^15] This pattern was also observed in robustness checks that used a flexible Bayesian Additive Regression Trees (BART) model for estimation. A state with Michigan’s dose (nearly 4) has an estimated CATE of around 1.25-1.25, with 95% credible interval (–2.0, –0.4), and so would be expected to gain just over 1 Democratic seat, on average. Similarly, a state with Ohio’s dose (just over 1) and covariates would be expected to gain about half of a Democratic seat (–0.6, –0.1). Motivated by this pattern, we also include in the right figure a dose-response curve calculated by calculating a CATE for dose level and averaging over the observed covariate distribution. [^16] Because of the interaction between the dose variable and the covariates (which do not have mean zero), the overall slope estimate and its variance differ from the estimate and variance of the coefficient on Dose in the linear model.
@@
In terms of partisan outcomes, the estimated ACRs on Republican seats are all positive, for both realized and maximum leeway. This is in line with an interpretation that increasing leeway benefits the party in control, as one might expect. Recall that the doses from Figure 4, for the realized leeway estimates, a dose of 4 represents approximately going from a Democratic legislature to an independent commission or an independent commission to a Republican legislature. For such a dose of 4, we would expect Republicans to gain around 0.7 seats (0.21, 1.18), or 7.2 simulation standard deviations (2.4, 12.2)—again, a substantial effect.
5.3 Placebo check
As a further validation of our overall approach to estimating causal effects, we conduct a placebo check by applying the same estimation procedure to a political outcome variable that should not be affected by redistricting reform but is closely related to the outcome variable of interest. For this placebo outcome variable, we use the Democratic two-party vote share in the first presidential election following each census (2012 and 2024). As national contests, presidential elections are insulated from state-specific reform efforts. Moreover, issues like the economy, immigration, and health care dominate presidential campaigns, not state-specific governance issues such as redistricting reform. Thus, we expect there to be no effect of redistricting reforms on presidential vote share.
Figure 8: Average causal response (ACR) of leeway on placebo outcome. The points correspond to the mean estimated ACR, while the lines represent 80% and 95% credible intervals. Positive numbers correspond to an increase in Democratic vote share.
Figure 8 displays the estimated ACRs for the placebo outcome using each of the three treatment variables. All three estimates are nearly zero, with credible intervals that easily cover zero. If the estimates were significantly different from zero, we would suspect the validity of the sufficiency assumption, the parallel trends assumption, or the estimation method. That all placebo estimates are close to zero is therefore strong evidence of the plausibility of these assumptions and our estimation strategy overall.
@@ 6 Redistricting Commissions: Counterfactual Policy Analysis
What if redistricting reforms were adopted nationwide? Redistricting commissions are the most commonly adopted map-drawing reform. A total of 15 states, including Michigan, New Jersey, and Arizona, currently use some form of commission. Advocates often argue that commissions can lead to fairer redistricting plans by further removing legislators from the process (e.g., American Academy of Arts & Sciences Commission on the Practice of Democratic Citizenship,, 2020). [^18] See also People Not Politicians Oregon or Ohio’s Citizens Not Politicians.
In this section, we conduct a counterfactual policy analysis of the nationwide adoption of redistricting commissions. Though hypothetical, counterfactual policy analyses are useful tools for evaluating institutional reform proposals (Cervas and Grofman,, 2019). Our approach is well suited for this analysis, in part because commissions can vary drastically in the extent to which they limit the influence of partisan actors.
6.1 Analysis procedure
To conduct the analysis, we apply fitted causal regression models based on the party-aware treatment to predict electoral outcomes under a series of hypothetical scenarios, in which every state adopts a redistricting commission with particular institutional powers, but partisan control of state institutions remains unchanged. Although each of the 15 states uses a redistricting commission with different structures and rules, our model is able to characterize these different commissions in terms of their partisan leeway and use these treatment values to predict electoral outcomes under a given commission structure.
We caution that any counterfactual policy analysis of this type involves substantial extrapolation from the observed data, and its results should therefore be interpreted as exploratory rather than confirmatory. While the regression models do allow for effect heterogeneity, the critical sufficiency assumption rules out unmodeled heterogeneity that may in practice lead to different reform effects.
@@ 6.1 Analysis procedure
We visualize the impact of each reform structure using a seats-votes curve. Seats-votes curves are often used to study a measure of partisan bias, operationalized as the difference in the share of seats and votes at a given point on the curve (Tufte,, 1973; King and Browning,, 1987; Katz et al.,, 2020). A seats-votes curve for which a 50% vote share translates exactly to a 50% seat share represents the baseline for partisan bias. We approximate the seats-votes curve around 50% vote share as a line, and estimate the effect of reforms on its slope and intercept separately. Up to a linear transformation, the slope corresponds to responsiveness and the intercept to seat share. These two outcome variables are differenced using the simulation sets, and a regression model is fit to these differenced outcomes, as discussed in Section 4.
6.2 Findings
[…]
@@ 6.2 Findings
We find that New York- and Ohio-style reforms produce relatively moderate improvements in partisan bias and responsiveness. The point estimates for the number of Democratic seats gained under these reforms is 3.3 for New York-style reforms, with 95% credible interval (0.75, 6.16), and 5.1 for Ohio-style reforms (1.3, 9.3). Responsiveness for New York- and Ohio-style reforms is also similar as for every 1pp increase in vote share, Democrats gain an estimated 8 seats with New York-style reforms, and 8.1 seats with Ohio-style reforms. These responsiveness values are only slightly higher than the actual responsiveness following 2020 redistricting. Ohio-style reforms appear more effective overall, though the substantive differences are small; we are more than 99% confident that Ohio-style reforms produce a larger effect on partisan bias than New York-style reforms and 88% confident that Ohio-style reforms increase responsiveness more than New York-style reforms do.
@@
Further, we find that Michigan-style reforms have the greatest effects in both responsiveness and Democratic seats, since partisan actors are the most constrained by the presence of a nonpartisan commission, no partisan veto points, and the potential for court review. We estimate that the number of Democratic seats gained under Michigan-style reforms is 6.2 with 95% credible interval (1.1, 11.9), and that for every 1pp increase in vote share, Democrats gain an estimated 8.1 seats (7.5, 8.7), an increase of 0.28 over the baseline. We are 83% confident that Michigan-style reforms have a larger effect than Ohio reforms, and 98% confident in a larger effect than New York-style reforms. We are also 82% confident that Michigan-style reforms increase responsiveness.
@@
All three proposed reforms reduce the deviation from partisan symmetry by increasing net Democratic seats. However, greater effects are produced when, as in Michigan-style reforms, multiple nodes of the redistricting game constrain leeway by ensuring partisan nodes are nonpartisan, and do not precede partisan vetoes. Meanwhile, examples of New York-style and Ohio-style reforms demonstrate how constraining partisan actors at different nodes, through different reforms, may produce substantively similar effects.